OpenStudy (highschoolmom2010):

Use the given circumference to find the surface area of each spherical object. Round your answer to the nearest whole number. 13. a grapefruit with C = 14 cm

OpenStudy (campbell_st):

well if the circumference is 14 cm use \[C = 2\pi r\] to find the radius so \[14 = 2\pi r\] just make r the subject. then substitute your value of r into the surface area of a sphere formula shown below \[SA = 4\pi r^2\] hope this helps

OpenStudy (highschoolmom2010):

to get r do i do 14/2 pi

OpenStudy (campbell_st):

thats right so its now \[SA = 4 \pi \times (\frac{14}{2\pi})^2\] now just do the arithmetic

OpenStudy (highschoolmom2010):

i must be doing something wrong because i got 1934 pi

OpenStudy (campbell_st):

probably the way you have entered it I've squared 14/2pi look at the common factors that cancel  now calculate